a company makes both LCD and CRT color monitors. the cost of materials for a CRT monitor is $50, whereas the cost of materials for a LCD monitor is $150. cost of labor to make a CRT monitor is $80, whereas the cost of labor to make an LCD monitor is $130. during a week when the company has budgeted $9600 for materials and $8760 for labor, how many LCD monitors does the company plan to make?
Answer by dukesmart(20)

let put x= CRT plan to be made
y= LCD plan to be made
we create the equations as follows:
50x+150y= 9600 ......(1)
80x+130y= 8760 ......(2)
equation (1) can be multiply by 8/5 will become; 80x + 240y = 15360 ...(3)
let take (3) subtract (2) we will have: 240y-130y=15360-8760
or: 110y= 6600 then y=60
plan to make 60 LCD and 12 CRT.